# Modal logics of almost sure validities in some classes of euclidean and transitive frames

### v1

@@ Abstract
We find complete and sound axiomatizations for the logics of almost sure validities in the classes of finite frames defined by the logics $\mathbf{K5}$, $\mathbf{KD5}$, $\mathbf{K45}$, $\mathbf{KD45}$, $\mathbf{K5B}$, $\mathbf{S5}$, $\mathbf{Grz.3}$, and $\mathbf{GL.3}.$
@@
Exploring the behavior of random structures for logical systems that extend beyond the first-order language is a compelling research direction. The modal language, interpreted on relational structures using Kripke semantics, is a particularly important example, which has led to several noteworthy discoveries. Le Bars disproved the zero-one law for modal logic in Kripke frames [Bar02] . Verbugge provided a valuable example of a modal logic that satisfies the zero-one law both in models and in frames, namely the provability logic GL, and gave countable axiomatizations for the logics of almost sure truths in this logic [Ver21].
@@
There are several results in the study of almost sure truths in the modal languages. The logic of almost sure truths in random Kripke models coincides with Carnap’s modal logic, according to [HK94]. The paper [GK03] offers a complete and sound axiomatization for the logic of almost sure validities in a countable frame. This result is utilized in [Gor20] to identify some of the almost sure validities in finite frames. However, the problem of complete axiomatization of almost sure validities in the class of finite frames remains open.
In this paper we discuss the logics of almost sure validities in various classes of finite frames. We generalize Goranko’s construction of the random finite frame: given a class of frames $\mathcal{C},$ we consider the uniform distribution on the labelled frames with $n$ states that belong to $\mathcal{C}.$ Our general result states that the almost sure validities form a normal modal logic that extends $\operatorname{Log}\mathcal{C}.$ We are most interested in the modally definable classes of frames. Using the tools of combinatorics and probability theory, we find finite axiomatizations for the logics of almost sure validities in the classes of frames defined by the logics $\mathbf{K5}$, $\mathbf{KD5}$, $\mathbf{K45}$, $\mathbf{KD45}$, $\mathbf{K5B}$, $\mathbf{S5}$, $\mathbf{Grz.3},$ and $\mathbf{GL.3}.$
@@ 1.1 Modal syntax and semantics
If $\varphi,\,\varphi\to\psi\in L$ then $\psi\in L,$
If $\phi\in L$ then $\Box\varphi\in L,$
@@
The notation $F\models\varphi$, where $F$ is a frame and $\varphi\in\mathrm{ML}$ is a formula, means ‘$\varphi$ is valid in $F$’ with the standard definition (see, for example, [BBW06]). For any set $\Gamma$ of modal formulas, $F\models\Gamma$ means $\forall\varphi\in\Gamma\>\left(F\models\varphi\right).$ The set $\operatorname{Log}F$ of all formulas $\varphi\in\mathrm{ML}$ that are valid in a frame $F$ is called the logic of $F.$ This definition extends to classes of frames in the usual way. Given a set of formulas $\Gamma\subseteq\mathrm{ML}$ and a number $n\in\omega,$ let $\operatorname{Fr}\Gamma$ denote the class of all frames $F$ such that $F\models\Gamma.$
@@ 1.2 Operations on frames
For any $U\subseteq X,$ $R[U]$ denotes the set $\{a\in X\mid\exists u\in U\>\left(uRa\right)\}.$
@@
If $F=(X,\,R)$ is a frame and $U\subseteq X$, the subframe of $X$ generated by $U$ is the frame $F{\uparrow}U=(R^{*}[U],\,R{\upharpoonright}R^{*}U).$ If $a\in X$ then $F{\uparrow}a$ is a shorthand for $F{\uparrow}\{a\}.$ A frame $F$ is said to be point-generated if $F=F{\uparrow}a$ for some $a\in\operatorname{dom}F.$
The generated subframe preserves the validity of modal formulas: for any $U\subseteq\operatorname{dom}F$, $\operatorname{Log}{F}\subseteq\operatorname{Log}{F{\uparrow}U}$ [BBW06].
@@
It is well-known that $\operatorname{Log}\biguplus_{i\in I}F_{i}=\bigcap_{i\in I}\operatorname{Log}F_{i}$ [BBW06].
Given a pair of frames $F=(X,\,R)$ and $G=(Y,\,S),$ a p-morphism from $F$ to $G$ is a surjective map $f:\>X\to Y$ such that $S(f(a))=f(R(a))$ for any $a\in X.$ We write $F\twoheadrightarrow G$ if there exists a p-morphism from $F$ to $G$.
The p-morphism preserves the validity of modal formulas: if $F\twoheadrightarrow G$ then $\operatorname{Log}F\subseteq\operatorname{Log}G.$
@@ 1.3 Classes of frames and their logics
In this paper we will consider the logics $\mathbf{K5},\,\mathbf{KD5},\,\mathbf{K45},\,\mathbf{KD45},\,\mathbf{K5B},\,\mathbf{S5},$ $\mathbf{GL.3},\,\mathbf{Grz.3}$. Recall that the frame classes of these logics are:
[…]
@@
These logics have the finite model property, so each of them is the logic of all finite point-generated frames that satisfy the corresponding frame condition [BdRV01]. For instance, $\mathbf{K5}$ is the logic of all finite point-generated Euclidean frames.
@@ 1.4 Random frames
For any $1\leq n\in\omega,$ let $\mathcal{F}_{n}=\{(n,\,R)\mid R\subseteq n\times n\}$ denote the set of all frames with the set of states $\{0,\,\ldots,\,n-1\}\equiv n.$
@@
For any property of frames $Q$, we say that $\hat{F}_{n}(\mathcal{C})$ has $Q$ asymptotically almost surely (a.a.s.) if the probability that $Q$ holds in $\hat{F}_{n}(C)$ tends to one as $n\to\infty:$
$\lim_{n\to\infty}\operatorname{P}(Q\text{ holds in }\hat{F}_{n}(\mathcal{C}))=1.$
@@
Finally, if $\varphi\in\operatorname{Log}(\mathcal{C})$ then $\varphi$ is valid in any possible value of $\hat{F}_{n}(\mathcal{C}).$ Then for any $n\in\omega,\,\operatorname{P}(\hat{F}_{n}(\mathcal{C})\models\varphi)=1$, so ${\varphi\in\operatorname{Log}^{\mathbf{as}}(\mathcal{C})}.$
@@
If $L_{1},\,L_{2}$ are modal logics, $L_{1}^{\mathbf{as}}=L_{2}$ then $L_{2}^{\mathbf{as}}=L_{2}.$
@@
By Theorem 1.1 $L_{1}\subseteq L_{1}^{\mathbf{as}}=L_{2},$ so $\operatorname{Fr}L_{2}\subseteq\operatorname{Fr}L_{1}$ and $\operatorname{Fr}L_{2}\cap\operatorname{Fr}L_{1}=\operatorname{Fr}L_{2}.$ Since $L_{1}^{\mathbf{as}}=L_{2},$
$\frac{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L_{2}\cap\operatorname{Fr}L_{1}\right|}{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L_{1}\right|}=\frac{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L_{2}\right|}{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L_{1}\right|}\to 1,\,n\to\infty.$
Then for any $\varphi\in L_{2}^{\mathbf{as}},$
$\operatorname{P}(\hat{F}_{n}(L_{1})\models\varphi)=\lim_{n\to\infty}\frac{\mathcal{F}_{n}\cap\operatorname{Fr}L_{1}\cap\operatorname{Fr}\{\varphi\}}{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L_{1}\right|}\geq\lim_{n\to\infty}\frac{\mathcal{F}_{n}\cap\operatorname{Fr}L_{2}\cap\operatorname{Fr}\{\varphi\}}{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L_{1}\right|}\\ =\lim_{n\to\infty}\frac{\mathcal{F}_{n}\cap\operatorname{Fr}L_{2}\cap\operatorname{Fr}\{\varphi\}}{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L_{1}\right|}\frac{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L_{1}\right|}{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L_{2}\right|}=\operatorname{P}(\hat{F}_{n}(L_{2})\models\varphi)\cdot 1=1.$
Therefore $L_{2}^{\mathbf{as}}\supseteq L_{1}^{\mathbf{as}}=L_{2}.$ By Theorem 1.1 $L_{2}\subseteq L_{2}^{\mathbf{as}}.$ Then $L_{2}^{\mathbf{as}}=L_{2}$.
@@ 1.5 Asymptotics
    $f=o(g),$ there exists $\alpha:\>\omega\to\mathbb{R}$ such that $f(n)=\alpha(n)g(n)$ for all $n\in\omega$ and $\lim_{n\to\infty}\alpha(n)=0;$
@@ 1.6 Set partitions and combinatorial numbers
The Bell number $B_{n}$ is defined as the number of distinct partitions of the set $n=\{0,\,\ldots,\,n-1\}$. Equivalently, $B_{n}$ is the number of distinct equivalence relations on $n.$
Given $r\in\omega,$ the number of partitions $\mathcal{U}$ of $n$ such that $|U|\leq r$ for any $U\in\mathcal{U}$ is denoted $G_{n,r}.$ The asymptotical and combinatorial behavior of $G_{n,r}$ is discussed in [MMW58].
@@
$G_{n,r}2^{n^{k}}=o(B_{n}),\,n\to\infty$ (3)
Given $n\in\omega$ and $m\leq n,$ the binomial coefficient $\binom{n}{m}$ is defined as the number of distinct $m$-element subsets of the set $n.$ The following estimation holds for any $n\in\omega$ and $m\leq n$ [SF14]:
@@
The binomial coefficient and the Bell number are connected by the identity [Lov93]
$\sum_{m=0}^{n}\binom{n}{m}B_{n}=B_{n+1},\quad n\in\omega.$
Since $B_{n}$ are increasing, it follows that $B_{n+1}\geq B_{n}\sum_{m=0}^{n}\binom{n}{m}=B_{n}2^{n},$ thus
$\frac{B_{n}}{B_{n+1}}=O(2^{-n})\to 0,\,n\to\infty.$ (5)
@@ 2 Connected frames
Let $n\in\omega$ and let $\mathcal{U}$ be a partition of $n.$ For any $U\in\mathcal{U}$ there exists a frame $F_{U}\in\mathcal{F}_{|U|}\cap\operatorname{Con}\operatorname{Fr}L.$ Construct a frame $F_{\mathcal{U}}=(n,R_{\mathcal{U}})$ by putting a copy of $F_{U}$ on $U$ for any $U\in\mathcal{U}.$ Then $F_{\mathcal{U}}$ is isomorphic to $\biguplus_{U\in\mathcal{U}}F_{U},$ and $F_{\mathcal{U}}\models L$ since $F_{U}\models L$ for any $U\in\mathcal{U}.$ Thus $F_{\mathcal{U}}\in\mathcal{F}_{n}\cap\operatorname{Fr}L.$
Observe that $\mathcal{U}$ is exactly the set of connected components of $F_{\mathcal{U}}.$ Thus if $\mathcal{U},\,\mathcal{V}$ are partitions of $X$ and $F_{\mathcal{U}}=F_{\mathcal{V}}$ then $\mathcal{U}=\mathcal{V}.$ Then $\mathcal{U}\mapsto F_{\mathcal{U}}$ is an injective mapping of the partitions of $n$ into $\mathcal{F}_{n}\cap\operatorname{Fr}L,$ so $|\mathcal{F}_{n}\cap\operatorname{Fr}L|\geq B_{n}.$
@@
Let $r\in\omega.$ Denote $\mathcal{F}_{n}^{\leq r}$ the set of frames in $\mathcal{F}_{n}$ whose connected components have cardinality at most $r.$ Then $\left|\mathcal{F}_{n}^{\leq r}\right|=o(B_{n}),\,n\to\infty.$
Proof 2.4.
Let $F=(n,R)\in\mathcal{F}_{n}^{\leq r}.$ If $\mathcal{U}$ is the set of connected components of $F$ then $(a,b)\not\in R$ for any $a\in U,\,b\in V$ where $U,\,V\in\mathcal{U},\,U\neq V.$ Then
@@
Then $R\in\mathcal{P}\left(\bigcup_{U\in\mathcal{U}}U\times U\right).$ By the hypothesis, $|U|\leq r$ for any $U\in\mathcal{U},$ so we can estimate
@@
$|\mathcal{F}_{n}^{\leq r}|\leq\sum_{\mathcal{U}}|A_{\mathcal{U}}|\leq\sum_{\mathcal{U}}2^{|\mathcal{U}|r^{2}}\leq\sum_{\mathcal{U}}2^{nr^{2}},$
where the sum is taken over all partitions $\mathcal{U}$ of $n$ into sets of cardinality at most $r.$ The number of such partitions is $G_{n,r},$ so by (3)
$|\mathcal{F}_{n}|^{\leq r}\leq G_{n,r}2^{nr^{2}}=o(G_{n,r}2^{n^{3}})=o(B_{n}).$
@@
Define a random subset $\hat{U}_{n}\subseteq n$ to be the connected component of $\hat{F}_{n}(L)$ that has the maximal cardinality. If there are more than one such components $U_{1},\,\ldots,\,U_{k}$, let $\hat{U}$ be the one that contains the state $a=\min\bigcup_{j=1}^{k}U_{k}$.
The generated subframe $\hat{F}_{n}(L){\uparrow}\hat{U}_{n}$ is a connected frame that validates $L$, so $\hat{F}_{n}(L){\uparrow}\hat{U}_{n}$ coincides, up to a monotone relabeling of states, with some frame from $\mathcal{F}_{|\hat{U}_{n}|}\cap\operatorname{Con}\operatorname{Fr}L.$
We claim that for any fixed $U\subseteq n$, the values of $\hat{F}_{n}(L){\uparrow}\hat{U}_{n}$ with $\hat{U}_{n}=U$ are distributed uniformly on $\mathcal{F}_{|U|}\cap\operatorname{Con}\operatorname{Fr}L.$ Informally, $\hat{F}_{n}(L){\uparrow}\hat{U}_{n}$ is independent of $\hat{F}_{n}(L){\uparrow}(n\setminus\hat{U}_{n})$.
Let $U\subseteq n$ and denote $\mathcal{G}_{n,U}$ the set of frames from $\mathcal{F}_{n}\cap\operatorname{Fr}L$ where $U$ is the maximal connected component. Let us consider any $F_{1},\,F_{2}\in\mathcal{F}_{|U|}\cap\operatorname{Con}\operatorname{Fr}L$. For any $G=(n,R)\in\mathcal{G}_{n,U}$ such that $G{\uparrow}U$ equals $F_{1}$ (up to a monotone relabeling, which we assume hereinafter), we construct a frame $G^{\prime}$ by changing the relation of $G$ on $U$ in such a way that $G^{\prime}{\uparrow}(n\setminus U)=G{\uparrow}(n\setminus U)$ and $G^{\prime}{\uparrow}U=F_{2}.$
Let us show that $G^{\prime}\in\mathcal{G}_{n,U}.$ By the construction, $U$ is the maximal connected component of $G^{\prime}.$ Moreover, $G\models L$ implies that $G{\uparrow}(n\setminus U)\models L,$ so $G^{\prime}\models L$ since $G^{\prime}\cong G{\uparrow}(n\setminus U)\,\uplus\,F_{2}$ and $F_{2}\models L.$ Thus $G^{\prime}\in\mathcal{F}_{n}\cap\operatorname{Fr}L.$
@@
consequently, for any $U\subseteq n$ and $F_{1},\,F_{2}\in\mathcal{F}_{|U|}\cap\operatorname{Con}\operatorname{Fr}L,$
@@
Since $\hat{F}_{n}(L){\uparrow}\hat{U}_{n}$ is a generated subframe of $\hat{F}_{n}(L),$ $\hat{F}_{n}(L)\not\models\varphi$ whenever $\hat{F}_{n}(L){\uparrow}\hat{U}_{n}\not\models\varphi.$ Then by the law of total probability
@@
$=\sum_{U\subseteq n}\operatorname{P}\left(\hat{F}_{n}(L){\uparrow}\hat{U}_{n}\not\models\varphi\,\Big{|}\,\hat{U}_{n}=U\right)P(\hat{U}_{n}=U)$

### v2

@@ Abstract
We find complete and sound axiomatizations for the logics of almost sure validities in the classes of finite frames defined by the logics $\mathbf{KD5}$, $\mathbf{KD45}$, $\mathbf{K5B}$, $\mathbf{S5}$, $\mathbf{Grz.3}$, and $\mathbf{GL.3}.$
Mathematocal Subject Classification: 03B45.
@@
Exploring the behavior of random structures for logical systems that extend beyond the first-order language is a compelling research direction. The modal language, interpreted on relational structures using Kripke semantics, is a particularly important example, which has led to several noteworthy discoveries. Le Bars disproved the zero-one law for modal logic in Kripke frames [LB02]. Verbrugge [Ver18] proved the zero-one law in the finite models of of Grzegorczyk’s logic and weak Grzegorczyk’s logic, and later provided a valuable example of a modal logic that satisfies the zero-one law both in models and in frames, namely the provability logic GL [Ver21].
@@
[…]
In this paper we discuss the logics of almost sure validities in various classes of finite frames. We generalize Goranko’s construction of the random finite frame: given a class of frames $\mathcal{C},$ we consider the uniform distribution on the labelled frames with $n$ states that belong to $\mathcal{C}.$ Our general result states that the almost sure validities form a normal modal logic that extends $\operatorname{Log}\mathcal{C}.$ We achieve an important technical result of studying such logics: for a large class of logics, the almost sure validities in a random $\mathcal{C}$-frame are also almost surely valid in a random connected $\mathcal{C}$-frame (Theorem 2.9). Since the connected frames typically have a simpler combinatorial structure, this connection allows us to find upper bounds on the logic of almost sure validities in $\mathcal{C}.$ We use this general theory to find finite axiomatizations for the logics of almost sure validities in the classes of frames defined by the logics $\mathbf{KD5}$, $\mathbf{KD45}$, $\mathbf{K5B}$, $\mathbf{S5}$, $\mathbf{Grz.3},$ and $\mathbf{GL.3}.$
@@ 1.1 Modal syntax and semantics
If $\varphi,\,\varphi\to\psi\in L$, then $\psi\in L,$
If $\phi\in L$, then $\Box\varphi\in L,$
@@
The notation $F\models\varphi$, where $F$ is a frame and $\varphi\in\mathrm{ML}$ is a formula, means ‘$\varphi$ is valid in $F$’ with the standard definition (see, for example, [BdRV01, Definition 1.28]). For any set $\Gamma$ of modal formulas, $F\models\Gamma$ means $\forall\varphi\in\Gamma\>\left(F\models\varphi\right).$ The set $\operatorname{Log}F$ of all formulas $\varphi\in\mathrm{ML}$ that are valid in a frame $F$ is called the logic of $F.$ This definition extends to classes of frames: if $\mathcal{F}$ is a class of frames, then $\operatorname{Log}\mathcal{F}$ is the set of all formulas that are valid in any frame $F\in\mathcal{F}$. Given a set of formulas $\Gamma\subseteq\mathrm{ML}$, let $\operatorname{Fr}\Gamma$ denote the class of all frames $F$ such that $F\models\Gamma.$
@@ 1.2 Operations on frames
For any $U\subseteq X,$ $R[U]$ denotes the set $\{a\in X\mid\exists u\in U\>\left(uRa\right)\}.$ The notation $R(a)$, where $a\in X,$ abbreviates $R[\{a\}].$
@@
If $F=(X,\,R)$ is a frame and $U\subseteq X$, the subframe of $X$ generated by $U$ is the frame $F{\uparrow}U=(R^{*}[U],\,R{\upharpoonright}R^{*}[U]).$ If $a\in X$, then $F{\uparrow}a$ is a shorthand for $F{\uparrow}\{a\}.$ A frame $F$ is said to be point-generated if $F=F{\uparrow}a$ for some $a\in\operatorname{dom}F.$ The generated subframe preserves the validity of modal formulas: for any $U\subseteq\operatorname{dom}F$, $\operatorname{Log}{F}\subseteq\operatorname{Log}{F{\uparrow}U}$ [BdRV01, Proposition 2.6].
@@
Given a pair of frames $F=(X,\,R)$ and $G=(Y,\,S),$ a p-morphism from $F$ to $G$ is a surjective map $f:\>X\to Y$ such that $S(f(a))=f(R(a))$ for any $a\in X.$ We write $F\twoheadrightarrow G$ if there exists a p-morphism from $F$ to $G$. The p-morphism preserves the validity of modal formulas: if $F\twoheadrightarrow G$, then $\operatorname{Log}F\subseteq\operatorname{Log}G$ [BdRV01, Proposition 2.14]
A frame isomorphism between $F=(X,\,R)$ and $G=(Y,\,S),$ is a bijection $f:\>X\to Y$ such that $aRb$ iff $f(a)Rf(b)$ for all $a,\,b\in X.$ It is straightforward that the existence of an isomorphism between $F$ and $G$ implies that $\operatorname{Log}F=\operatorname{Log}G.$
@@ 1.3 Classes of frames and their logics
In this paper we will consider the logics $\mathbf{KD5},\,\mathbf{KD45},\,\mathbf{K5B},\,\mathbf{S5},$ $\mathbf{GL.3},\,\mathbf{Grz.3}$. Recall that the frame classes of these logics are:
@@
These logics have the finite model property, so each of them is the logic of all finite point-generated frames that satisfy the corresponding frame condition [BdRV01]. For instance, $\mathbf{KD5}$ is the logic of all finite point-generated serial Euclidean frames.
@@ 1.4 Random frames
For any $1\leq n\in\omega,$ let $[n]$ denote the set $\{0,\,\ldots,\,n-1\}$, and let $\mathcal{F}_{n}=\{([n],\,R)\mid R\subseteq[n]\times[n]\}$ be the set of all frames with the set of states $[n].$
@@
Formally, we fix some measure space $(\Omega,\,\mathcal{G},\,\operatorname{P}),$ where $\mathcal{G}$ is a $\sigma$-algebra on $\Omega$ and $\operatorname{P}:\>\mathcal{G}\to[0,1]$ is a measure, and define $\hat{F}_{n}(\mathcal{C})$ to be a measurable map from $\Omega$ to $\mathcal{F}_{n}$ that satisfies (1), where $\operatorname{P}(\hat{F}_{n}(\mathcal{C})\in A)$ is a notation for  $\operatorname{P}\{\omega\in\Omega\mid\hat{F}_{n}(\mathcal{C})(\omega)\in A\}.$ The values $\hat{F}_{n}(\mathcal{C})(\omega)$ for $\omega\in\Omega$ are called realizations of $\hat{F}_{n}(\mathcal{C}).$
For any set of frames $\mathcal{Q}\subseteq\mathcal{F}_{n}$, we say that $\hat{F}_{n}(\mathcal{C})$ belongs to $\mathcal{Q}$ asymptotically almost surely (a.a.s.) if
$\lim_{n\to\infty}\operatorname{P}(\hat{F}_{n}(\mathcal{C})\in Q)=1.$
Sometimes we will refer to a set of frames $\mathcal{Q}\subseteq\mathcal{F}_{n}$ as a property of frames. In this case ‘$\mathcal{Q}$ holds in $\hat{F}_{n}(\mathcal{C})$ a.a.s.’ means that $\hat{F}_{n}(\mathcal{C})\in\mathcal{Q}$ a.a.s.
@@
Finally, if $\varphi\in\operatorname{Log}(\mathcal{C})$, then $\varphi$ is valid in any possible value of $\hat{F}_{n}(\mathcal{C}).$ Then for any $n\in\omega,\,\operatorname{P}(\hat{F}_{n}(\mathcal{C})\models\varphi)=1$, so ${\varphi\in\operatorname{Log}^{\mathbf{as}}(\mathcal{C})}.$
@@
If $L$ is a modal logic, then $(L^{\mathbf{as}})^{\mathbf{as}}=L^{\mathbf{as}}.$
@@
By Theorem 1.1 $L^{\mathbf{as}}\subseteq(L^{\mathbf{as}})^{\mathbf{as}}.$ For the other direction, observe that $\operatorname{Fr}L^{\mathbf{as}}\subseteq\operatorname{Fr}L$ since $L\subseteq L^{\mathbf{as}}$, so for any $\varphi\in(L^{\mathbf{as}})^{\mathbf{as}}$,
$\operatorname{P}(\hat{F}_{n}(L)\models\varphi)$ $=\lim_{n\to\infty}\frac{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L\cap\operatorname{Fr}\{\varphi\}\right|}{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L\right|}$
$\geq\lim_{n\to\infty}\frac{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L^{\mathbf{as}}\cap\operatorname{Fr}\{\varphi\}\right|}{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L\right|}$
$=\lim_{n\to\infty}\frac{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L^{\mathbf{as}}\cap\operatorname{Fr}\{\varphi\}\right|}{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L^{\mathbf{as}}\right|}\cdot\frac{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L^{\mathbf{as}}\right|}{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L\right|}$
$=\lim_{n\to\infty}\frac{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L^{\mathbf{as}}\cap\operatorname{Fr}\{\varphi\}\right|}{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L^{\mathbf{as}}\right|}\lim_{n\to\infty}\frac{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L^{\mathbf{as}}\right|}{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L\right|}.$
$=\operatorname{P}(\hat{F}_{n}(L^{\mathbf{as}})\models\varphi)\cdot\operatorname{P}(\hat{F}_{n}(L)\in\operatorname{Fr}L^{\mathbf{as}})=1\cdot 1=1.$
@@ 1.5 Asymptotics
    $f=o(g),$ if there exists $\alpha:\>\omega\to\mathbb{R}$ such that $f(n)=\alpha(n)g(n)$ for all $n\in\omega$ and $\lim_{n\to\infty}\alpha(n)=0;$
@@ 1.6 Set partitions and combinatorial numbers
The Bell number $B_{n}$ is defined as the number of distinct partitions of the set $[n]=\{0,\,\ldots,\,n-1\}$. Equivalently, $B_{n}$ is the number of distinct equivalence relations on $[n].$ The growth rate of the Bell number is described by the asymptotic expression [dB58, Section 6.2]
$\ln B_{n}=n(\ln n-\ln\ln n-1+o(1)),\quad n\to\infty.$ (3)
The Bell numbers satisfy
$\lim_{n\to\infty}\frac{B_{n}}{B_{n+1}}\to 0.$ (4)
We give the proof of (4) in Appendix.
Given $r\in\omega,$ the number of partitions $\mathcal{U}$ of $[n]$ such that $|U|\leq r$ for any $U\in\mathcal{U}$ is denoted $G_{n,r}.$ The asymptotic and combinatorial behavior of $G_{n,r}$ is discussed in [MMW58].
@@
$G_{n,r}2^{kn}=o(B_{n}),\,n\to\infty$ (5)
Given $n\in\omega$ and $m\leq n,$ the binomial coefficient $\binom{n}{m}$ is defined as the number of distinct $m$-element subsets of the set $[n].$ The following estimation holds for any $n\in\omega$ and $m\leq n$ [SF14, Section 5.4]:
@@ 2 Connected frames
Let $n\in\omega$ and let $\mathcal{U}$ be a partition of $[n].$ For any $U\in\mathcal{U}$ there exists a frame $F_{U}\in\mathcal{F}_{|U|}\cap\operatorname{Con}\operatorname{Fr}L.$ Construct a frame $F_{\mathcal{U}}=([n],R_{\mathcal{U}})$ by putting a copy of $F_{U}$ on $U$ for any $U\in\mathcal{U}.$ Then $F_{\mathcal{U}}$ is isomorphic to $\biguplus_{U\in\mathcal{U}}F_{U},$ and $F_{\mathcal{U}}\models L$ since $F_{U}\models L$ for any $U\in\mathcal{U}.$ Thus $F_{\mathcal{U}}\in\mathcal{F}_{n}\cap\operatorname{Fr}L.$
Observe that $\mathcal{U}$ is exactly the set of connected components of $F_{\mathcal{U}}.$ Thus if $\mathcal{U},\,\mathcal{V}$ are partitions of $X$ and $F_{\mathcal{U}}=F_{\mathcal{V}}$, then $\mathcal{U}=\mathcal{V}.$ Then $\mathcal{U}\mapsto F_{\mathcal{U}}$ is an injective mapping of the partitions of $[n]$ into $\mathcal{F}_{n}\cap\operatorname{Fr}L,$ so $|\mathcal{F}_{n}\cap\operatorname{Fr}L|\geq B_{n}.$
Definition 2.3.
Let $r\in\omega.$ Denote by $\mathcal{F}_{n}^{\leq r}$ the set of frames in $\mathcal{F}_{n}$ whose connected components have cardinality at most $r.$
@@
For any $r\in\omega,$ $\left|\mathcal{F}_{n}^{\leq r}\right|=o(B_{n}),\,n\to\infty.$
Proof 2.5.
Let $F=([n],R)\in\mathcal{F}_{n}^{\leq r}.$ If $\mathcal{U}$ is the set of connected components of $F$, then $(a,b)\not\in R$ for any $a\in U,\,b\in V$ where $U,\,V\in\mathcal{U},\,U\neq V.$ Then
@@
Then $R\in\mathcal{P}\left(\bigcup_{U\in\mathcal{U}}U\times U\right).$ By the assumption, $|U|\leq r$ for any $U\in\mathcal{U},$ so we can estimate
@@
$|\mathcal{F}_{n}^{\leq r}|\leq\sum_{\mathcal{U}}|A_{\mathcal{U}}|\leq\sum_{\mathcal{U}}2^{r^{2}|\mathcal{U}|}\leq\sum_{\mathcal{U}}2^{r^{2}n},$
where the sum is taken over all partitions $\mathcal{U}$ of $[n]$ into sets of cardinality at most $r.$ The number of such partitions is $G_{n,r},$ so by (5)
$|\mathcal{F}_{n}|^{\leq r}\leq G_{n,r}2^{r^{2}n}=o(B_{n}).$
@@
To prove the following theorem, we consider the distributions of generated subframes of the random frame that have some fixed size $m\in\omega.$ To simplify the reasoning, we view them as random frames in $\mathcal{F}_{m},$ using the following definition.
Definition 2.8.
Let $m,\,n\in\omega,\,m\leq n.$ If $U\subseteq[n]$ and $|U|=m,$ the monotone relabeling of $U$ is the unique bijection $\alpha:\>U\to[m]$ such that $\alpha(a)\leq\alpha(b)$ iff $a\leq b.$ Two frames $F=(U,R)$ and $G=([m],S)$ coincide up to monotone relabeling if the monotone relabeling $\alpha:\>U\to[m]$ is a frame isomorphism between $F$ and $G.$
@@
Define a random subset $\hat{U}_{n}\subseteq[n]$ to be the connected component of $\hat{F}_{n}(L)$ that has the maximal cardinality. If there are more than one such components $U_{1},\,\ldots,\,U_{k}$, let $\hat{U}$ be the one that contains the state $a=\min\bigcup_{j=1}^{k}U_{k}$.
Any realization of the generated subframe $\hat{F}_{n}(L){\uparrow}\hat{U}_{n}$ is a connected frame that validates $L$, so $\hat{F}_{n}(L){\uparrow}\hat{U}_{n}$ coincides, up to monotone relabeling of states, with some frame from $\mathcal{F}_{|\hat{U}_{n}|}\cap\operatorname{Con}\operatorname{Fr}L$ almost surely.
We claim that for any fixed $U\subseteq[n]$, the values of $\hat{F}_{n}(L){\uparrow}\hat{U}_{n}$ with $\hat{U}_{n}=U$ are distributed uniformly on $\mathcal{F}_{|U|}\cap\operatorname{Con}\operatorname{Fr}L.$ Informally, $\hat{F}_{n}(L){\uparrow}\hat{U}_{n}$ is independent of $\hat{F}_{n}(L){\uparrow}([n]\setminus\hat{U}_{n})$.
Let $U\subseteq[n]$ and denote by $\mathcal{G}_{n,U}$ the set of frames from $\mathcal{F}_{n}\cap\operatorname{Fr}L$ where $U$ is the maximal connected component. Let us consider any $F_{1},\,F_{2}\in\mathcal{F}_{|U|}\cap\operatorname{Con}\operatorname{Fr}L$. For any $G=([n],R)\in\mathcal{G}_{n,U}$ such that $G{\uparrow}U$ equals $F_{1}$ (up to monotone relabeling, which we assume hereinafter), we construct a frame $G^{\prime}$ by changing the relation of $G$ on $U$ in such a way that $G^{\prime}{\uparrow}([n]\setminus U)=G{\uparrow}([n]\setminus U)$ and $G^{\prime}{\uparrow}U=F_{2}.$
Let us show that $G^{\prime}\in\mathcal{G}_{n,U}.$ By the construction, $U$ is the maximal connected component of $G^{\prime}.$ Moreover, $G\models L$ implies that $G{\uparrow}([n]\setminus U)\models L,$ so $G^{\prime}\models L$ since $G^{\prime}\cong G{\uparrow}([n]\setminus U)\,\uplus\,F_{2}$ and $F_{2}\models L.$ Thus $G^{\prime}\in\mathcal{F}_{n}\cap\operatorname{Fr}L.$
@@
consequently, for any $U\subseteq[n]$ and $F_{1},\,F_{2}\in\mathcal{F}_{|U|}\cap\operatorname{Con}\operatorname{Fr}L,$
@@
Since $\hat{F}_{n}(L){\uparrow}\hat{U}_{n}$ is a generated subframe of $\hat{F}_{n}(L),$ $\hat{F}_{n}(L)\not\models\varphi$ whenever $\hat{F}_{n}(L){\uparrow}\hat{U}_{n}\not\models\varphi.$ Then by the law of total probability [Gut13, Proposition 4.1]
@@
$=\sum_{U\subseteq[n]}\operatorname{P}\left(\hat{F}_{n}(L){\uparrow}\hat{U}_{n}\not\models\varphi\,\Big{|}\,\hat{U}_{n}=U\right)P(\hat{U}_{n}=U)$